# 1FLAT: a Firmamento-based catalog of AGN in Fermi-LAT high Galactic latitude $\gamma$-ray sources

### v1

@@ 4.2 Synchrotron peak frequency and blazar classes
Table 4: Classification summary for blazars. ”Same” are those that both LAT and [Uncaptioned image] finds. ”New” are those found only in [Uncaptioned image]. ”Different” are the alternative blazar [Uncaptioned image] founds with respect to LAT.
@@ 4.3 Spectral properties
To further compare the sample of new associations with the confirmed associations in 4FGL in Fig. 9 we plot the histogram of $\gamma$-ray  spectral indices in the subsamples of HSP and LSP blazars, for the large sample of confirmed blazars and for the newly discovered blazars, along with some statistics. The distributions of $\gamma$-ray  spectral indices of the new blazars appear similar to that of the confirmed blazars. Therefore we conclude that our new associations are an extension of the same type of blazars found in 4FGL rather than a new population with peculiar properties.
@@ 6 Discussion and Conclusions
In this study, we have presented the 1FLAT , an independently derived catalog of blazar counterparts of high-Galactic latitude $\gamma$-ray sources in the LAT 4FGL catalog, utilizing the [Uncaptioned image] web-based platform. Our results showed a high level of agreement (over 88%) with the associations reported in the 4FGL and 4LAC catalogs, but also significant differences. A key finding of this work is the identification of 421 new credible blazar associations for previously unassociated $\gamma$-ray sources in the 4FGL catalog. This significantly reduces the fraction of unidentified high-Galactic latitude LAT sources from approximately 25% to 17%. We also identified 64 alternative blazar counterparts and found 49 cases where the 4FGL/4LAC association was not confirmed.
@@
Future studies could address these limitations and further enhance the 1FLAT  catalog. Incorporating quantitative variability measures more systematically and developing more sophisticated machine learning techniques for automated SED classification and counterpart association could improve the robustness and efficiency of the process. The upcoming release of the full eROSITA X-ray survey data is expected to significantly aid in identifying counterparts for currently unassociated $\gamma$-ray sources, providing an opportunity to further validate and expand the 1FLAT  catalog.
@@
In summary, the 1FLAT catalog, available online through the [Uncaptioned image] platform, offers a complementary and independently derived list of AGN associated with LAT $\gamma$-ray sources, and is intended as a useful reference for very high energy $\gamma$-ray observations, multi-wavelength campaigns, and studies of $\gamma$-ray emitting AGN, particularly in light of the significant number of new associations presented here.
@@ Appendix B Step by step procedure for source association
Fermi-LAT 4FGL-DR4 related entries $\downarrow$
@@
Fermi-LAT 4LAC-DR3 related entries $\downarrow$
@@
BLAZAR confirmed For blazars in which [Uncaptioned image] agrees with Fermi-LAT in unassisted way
BLAZAR confirmed visual For blazars in which [Uncaptioned image] agrees with Fermi-LAT in assisted way
@@
UNCERTAIN For sources in which [Uncaptioned image] does not find a valid candidate whereas Fermi-LAT does
NOC confirmed For sources in which [Uncaptioned image] agrees with Fermi-LAT that no counterparts are found
NOC new For sources where [Uncaptioned image] does not find a counterpart whereas Fermi-LAT does
GALAXY confirmed For galaxies in which [Uncaptioned image] agrees with Fermi-LAT
GALAXY new For galaxies in which [Uncaptioned image] disagrees with Fermi-LAT

### v2

@@ 4.2 Synchrotron peak frequency and blazar classes
Table 4: Classification summary for blazars. ”Same” are those that both 4FGL and [Uncaptioned image] find. ”New” are those found only in [Uncaptioned image]. ”Different” are the alternative blazar [Uncaptioned image] finds with respect to 4FGL.
@@ 4.3 Spectral properties
To further compare the sample of new associations with the confirmed associations in 4FGL, we plot in Fig. 9 the histogram of $\gamma$-ray  spectral indices for the HSP and LSP blazar subsamples. The histograms show both the large sample of confirmed blazars and the newly discovered blazars, along with some relevant statistics. The distributions of $\gamma$-ray  spectral indices of the new blazars appear similar to that of the confirmed blazars. Therefore we conclude that our new associations are an extension of the same type of blazars found in 4FGL rather than a new population with peculiar properties.
@@ 6 Discussion and Conclusions
In this study, we have presented the 1FLAT , an independently derived catalog of blazar counterparts of high-Galactic latitude $\gamma$-ray  sources in the LAT 4FGL catalog, constructed using the [Uncaptioned image]  web-based platform. Our results showed a high level of agreement (over 88%) with the associations reported in the 4FGL and 4LAC catalogs, but also significant differences. A key finding of this work is the identification of 421 new credible blazar associations for previously unassociated $\gamma$-ray sources in the 4FGL catalog. This significantly reduces the fraction of unidentified high-Galactic latitude LAT sources from approximately 25% to 17%. We also identified 64 alternative blazar counterparts and found 49 cases where the 4FGL/4LAC association was not confirmed.
@@
Future studies could address these limitations and further enhance the 1FLAT  catalog. Incorporating quantitative variability measures in a more systematic way could improve the robustness of the process. In addition, developing more sophisticated machine-learning techniques for automated SED classification and counterpart association would further enhance its efficiency. The upcoming release of the full eROSITA X-ray survey data is expected to significantly aid in identifying counterparts for currently unassociated $\gamma$-ray sources, providing an opportunity to further validate and expand the 1FLAT  catalog.
@@
In summary, the 1FLATcatalog, available online through the [Uncaptioned image]  platform, provides a complementary and independently derived list of AGN associated with LAT $\gamma$-ray  sources. It is intended as a useful reference for very-high-energy $\gamma$-ray  observations, multi-wavelength campaigns, and studies of $\gamma$-ray–emitting AGN, particularly in view of the significant number of new associations presented here.
@@ Appendix B Step by step procedure for source association
LAT 4FGL-DR4 related entries $\downarrow$
@@
LAT 4LAC-DR3 related entries $\downarrow$
@@
BLAZAR confirmed For blazars in which [Uncaptioned image] agrees with LAT in unassisted way
BLAZAR confirmed visual For blazars in which [Uncaptioned image] agrees with LAT in assisted way
@@
UNCERTAIN For sources in which [Uncaptioned image] does not find a valid candidate whereas LAT does
NOC confirmed For sources in which [Uncaptioned image] agrees with LAT that no counterparts are found
NOC new For sources where [Uncaptioned image] does not find a counterpart whereas LAT does
GALAXY confirmed For galaxies in which [Uncaptioned image] agrees with LAT
GALAXY new For galaxies in which [Uncaptioned image] disagrees with LAT